OpenStudy (anonymous):

Find the values of the six trigonometric functions of theta Function value Constraint sin theta= 3/5 theta lies in quadrant 2 PLEASE TELL ME HOW TEST TOMORROW

OpenStudy (mertsj):



OpenStudy (mertsj):

cos = -4/5 tan=-3/4 cot=-4/3 sec=-5/4 csc=5/3

OpenStudy (anonymous):

How do i know like opp hyp and adjacent :[ like how to draw the triangle and why cant the triangle be facing toward the center like flipped ?

OpenStudy (mertsj):

Because the vertex must be at the origin and the problem said it is in the second quadrant.

OpenStudy (mertsj):

Do you know the definitions of the functions?

OpenStudy (anonymous):

Soo it will always be like that like if its in quadrant 4 and 3 it will look like and i know soh cah toa I dont know cot csc and sec or maybe i do somthing is sin/cos i think 1/tan idk though

OpenStudy (mertsj):

yep